Flourine reacts with uranium to produce uranium hexafluoride,UF6. how mary fluorine molecules are required to produce 2.0mg of uranium hexafluoride from an excess of uranium? the molar mass of UF6 is 352gm/mol. 

Grade:12

2 Answers

Adarsh
8 years ago
U+3F2  → UF6
=>  3 x 6.02 x 1023 molecule F2 gives 352g UF6
=>  6.02 x 1023molecule F2 gives 352/3g UF6
=> 1g UF6 formed from (3/352)x 6.02 x 1023 F2 molecules.
=> 2mg UF6 formed from 2 x10-3 (3/352) x 6.02 x 1023 F2 molecules.
=> 2mg UF6 formed from1.026 x 1019 F2 molecules.
Please approve if my answer is correct.
 
Pallak
3 years ago
U(s)+3F6---->UF6
Number of moles of UF6=  2÷(1000 x 352)=5.68 x 10-6
1 mole of UF6=3 moles of F2
2÷(1000 x 352) moles of UF6=3 x 5.68 x 10-6 moles of F2
number of molecules of F2= 6.023 x 1023 x 3 x 5.68 x 10-6 = 1.026 x 1019 = 1 x 1019(approx)